# Score-based Neural Ordinary Differential Equations for Computing Mean Field Control Problems

### v1

@@ A.1 Proofs for score-based normalizing flow
If we compute the derivative of the MA equation equation 1 w.r.t. $t$, we obtain
@@
$=\left[\partial_{t}\rho(t,T(t,x))+\nabla_{T}\rho(t,T(t,x))^{\top}f(t,T(t,x))+\rho(t,T(t,x))\nabla_{T}\cdot f(t,T(t,x))\right]$
$\quad\cdot\det\left(\nabla_{x}T(t,x)\right)$
@@
where we have used equation 35 in the fourth equality. Since $T(t,\cdot)$ is invertible, $\det\left(\nabla_{x}T(t,x)\right)\neq 0$, which implies
@@
$\partial_{t}z_{t}$ $=f(t,z_{t})\,,$ (37)
$\partial_{t}l_{t}$ $=-\nabla_{z}\cdot f(t,z_{t})\,,$ (38)
$\partial_{t}s_{t}$ $=-\nabla_{z}f(t,z_{t})^{\top}s_{t}-\nabla_{z}(\nabla_{z}\cdot f(t,z_{t}))\,,$ (39)
$\partial_{t}H_{t}$ $=-\sum_{i=1}^{d}s_{it}\nabla_{z}^{2}f_{i}(t,z_{t})-\nabla_{z}^{2}(\nabla_{z}\cdot f(t,z_{t}))-H_{t}\nabla_{z}f(t,z_{t})-\nabla_{z}f(t,z_{t})^{\top}H_{t}\,,$ (40)
@@
The dynamic for $z_{t}$ equation 3a comes from the definition directly:
@@
We next show equation 4 and equation 3b.
@@
where we have used the continuity equation equation 2 in the fourth equality. Consequently,
@@
We next show equation 3c. We will use the fact frequently that $\nabla_{z}\log\rho(t,z_{t})=\nabla_{z}\rho(t,z_{t})/\rho(t,z_{t})$. We first compute $\dfrac{\mathrm{d}}{\mathrm{d}t}(\nabla_{z}\rho(t,z_{t}))$
$\quad\dfrac{\mathrm{d}}{\mathrm{d}t}\left[\nabla_{z}\rho(t,z_{t})\right]=(\partial_{t}\nabla_{z})\rho(t,z_{t})+\nabla_{z}^{2}\rho(t,z_{t})\,\partial_{t}z_{t}=\nabla_{z}\partial_{t}\rho(t,z_{t})+\nabla_{z}^{2}\rho(t,z_{t})\,f(t,z_{t})$ (41)
@@
where we have used equation 41 and equation 4 in the third inequality.
Finally we prove equation 3d. In order to simplify notation, we will omit $(t,z_{t})$ when there is no confusion and rewrite $\rho(t,z_{t})$ and $f(t,z_{t})$ as $\rho$ and $f$. We also denote $\partial_{i}$ as the partial derivative w.r.t. the $i$-th variable in $z$. We first compute $\dfrac{\mathrm{d}}{\mathrm{d}t}\nabla_{z}^{2}\rho(t,z_{t})$.
@@
$=-\nabla_{z}^{2}[\nabla_{z}\rho^{\top}f+\rho\,\nabla_{z}\cdot f]+\nabla_{z}^{3}\rho\cdot f=-\nabla_{z}^{2}\left[\sum_{i=1}^{d}\partial_{i}\rho\,f_{i}+\rho\,\nabla_{z}\cdot f\right]+\nabla_{z}^{3}\rho\cdot f$
$=-\sum_{i=1}^{d}\left(\nabla_{z}^{2}\partial_{i}\rho\,f_{i}+\partial_{i}\rho\,\nabla_{z}^{2}f_{i}+\nabla_{z}\partial_{i}\rho\,\nabla_{z}f_{i}^{\top}+\nabla_{z}f_{i}\,\nabla_{z}\partial_{i}\rho^{\top}\right)$
$\quad-\left(\nabla_{z}^{2}\rho\,(\nabla_{z}\cdot f)+\rho\,\nabla_{z}^{2}(\nabla_{z}\cdot f)+\nabla_{z}\rho\,\nabla_{z}(\nabla_{z}\cdot f)^{\top}+\nabla_{z}(\nabla_{z}\cdot f)\,\nabla_{z}\rho^{\top}\right)+\nabla_{z}^{3}\rho\cdot f$
@@
where we have used equation equation 44 in the second equality. Finally, we get
@@
We have used equation 43 in the second equality and used equation 45, equation 44, equation 4, and equation 41 in the fourth equality. This finishes the proof. ∎
@@
if equation equation 47 holds. We first compute
$\dfrac{\mathrm{d}}{\mathrm{d}t}\det(\Sigma(t))$ $=\det(\Sigma(t))\operatorname{Tr}\left(\Sigma(t)^{-1}\,\partial_{t}\Sigma(t)\right)=\det(\Sigma(t))\operatorname{Tr}\left(\Sigma(t)^{-1}\,(A(t)\Sigma(t)+\Sigma(t)A(t)^{\top})\right)$ (49)
$=\det(\Sigma(t))\operatorname{Tr}\left(\Sigma(t)^{-1}A(t)\Sigma(t)+A(t)^{\top}\right)=2\det(\Sigma(t))\operatorname{Tr}(A(t))\,,$ (50)
@@
$\dfrac{\mathrm{d}}{\mathrm{d}t}\left(\det(\Sigma(t))^{-\frac{1}{2}}\right)=-\frac{1}{2}\det(\Sigma(t))^{-\frac{3}{2}}\,\dfrac{\mathrm{d}}{\mathrm{d}t}\det(\Sigma(t))=-\det(\Sigma(t))^{-\frac{1}{2}}\,\operatorname{Tr}(A(t))\,.$ (51)
@@
Combining equation 51 and equation 53, we obtain
@@ A.2 Details for the MFC problems
$\inf_{v}\int_{0}^{t_{\text{end}}}\int_{\mathbb{R}^{d}}\left[L(t,x,v(t,x))+F(t,x,\rho(t,x))\right]\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t+\int_{\mathbb{R}^{d}}G(x,\rho(t_{\text{end}},x))\rho(t_{\text{end}},x)\,\mathrm{d}x\,,$ (56)
@@
Let $L$ be strongly convex in $v$, then solution to the MFC problem equation 23 is as follows. Consider a function $\psi\colon[0,t_{\textrm{end}}]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that
@@
$\left.\quad+\phi(t,x)\left(\partial_{t}\rho(t,x)+\nabla_{x}\cdot(f(t,x)\rho(t,x))\right)\right]\,\mathrm{d}x\,\mathrm{d}t+\int_{\mathbb{R}^{d}}G(x,\rho(t_{\text{end}},x))\rho(t_{\text{end}},x)\,\mathrm{d}x$ (66)
@@
Taking the variation of equation 68 w.r.t. $\rho(t_{\text{end}},\cdot)$, we get $\phi(t_{\text{end}},x)=-\widetilde{G}(x,\rho(t_{\text{end}},x))$. Taking the variation w.r.t. $f$, we obtain
@@
due to equation 60. Next, we denote $\mathrm{D}_{p}H(t,x,\nabla_{x}\phi(t,x))$ by $\mathrm{D}_{p}H$ for short. We then observe that
@@
Taking the variation of equation 68 w.r.t. $\rho(\cdot,\cdot)$, we get
@@
$=-\gamma\Delta_{x}\phi(t,x)+L(t,x,\mathrm{D}_{p}H)+\widetilde{F}(t,x,\rho(t,x))-\partial_{t}\phi(t,x)-\nabla_{x}\phi(t,x)^{\top}\mathrm{D}_{p}H\,,$ (75)
where we have used $\mathrm{D}_{v}L=\nabla_{x}\phi(t,x)$ in the last inequality. Equation equation 75 is the classical HJB equation corresponding to the MFC problem, which coincide with equation (4.6) in Bensoussan et al. (2013) after a sign flip, where our $\mathrm{D}_{p}H$ is their $\hat{v}$. Please note that, throughout our derivation, $\rho(t,x)$ is not a general density, but the density under the optimal control.
@@
where the last equality is due to equation 70. Therefore, the HJB equation equation 75 is transformed into
@@
Here, $\mathrm{D}_{p}H$ is short for $\mathrm{D}_{p}H(t,x,\nabla_{x}\psi(t,x)+\gamma\nabla_{x}\log\rho(t,x))$. We plug in the shift equation 76 in the first equality; we plug in equation 77 in the second equality. Therefore, we have recovered Proposition 3. ∎
@@
In the LQ problem where $L(t,x,v)=\frac{1}{2}|v|^{2}$ and $F(t,x,\rho)=0$, the solution of the MFC problem equation 23 is characterized by
@@ A.3 Formulation for flow matching of overdamped Langevin dynamic
For flow matching problem of overdamped Langevin dynamic where $L(t,x,v)=\frac{1}{2}|v-b(t,x)|^{2}$, $F(t,x,\rho)=0$ and $G(x,\rho)=0$, the solution of the MFC problem equation 23 is characterized by
@@ B.1 HJB regularizer for MFC
In addition to the numerical algorithm in section 4.3, we can add the residual of the HJB equation equation 27 into the loss function. First, let us consider the HJB regularizer in the LQ Gaussian case, as presented in Corollary 1. Here the Gaussian case means that $\rho_{0}$ is also given as a Gaussian distribution, in which the solution of density function in LQ MFC problem stays in a time dependent Gaussian distribution.
In the LQ scenario, parametrizing the vector field $f$ directly as a neural network complicates the system, as terms like $\partial_{t}\psi$ in equation 29 are hard to implement. A more feasible approach is to parametrize $\psi$ as a neural network, allowing the vector field $f$ to be computed as $f(t,x)=\nabla_{x}\psi(t,x)$.
This parametrization, however, requires two additional orders of auto-differentiations compared to the original algorithm. First, if $\psi$ is parametrized as a neural network as in equation 8, then $f(t,x)=\nabla_{x}\psi(t,x)$ must be obtained through auto-differentiation. Second, the term $\Delta_{x}\rho(t,x)$ in equation 29 requires computing the dynamic for $H_{t}$ through equation 13, which requires another order of derivative. This term $H_{t}$ is not required in the original Algorithm 1. These two additional differentiations could potentially complicate the optimization landscape.
@@
where $A(t):[0,t_{\text{end}}]\to\mathbb{R}^{d\times d}$, $B(t):[0,t_{\text{end}}]\to\mathbb{R}^{d}$, and $C(t):[0,t_{\text{end}}]\to\mathbb{R}$. These three functions can be further parametrized into trainable variables $\theta=[\theta^{A},\theta^{B},\theta^{C}]$ with $\theta^{A}\in\mathbb{R}^{(N_{t}+1)\times d\times d}_{\text{sym}}$, $\theta^{B}\in\mathbb{R}^{(N_{t}+1)\times d}$, and $\theta^{C}\in\mathbb{R}^{N_{t}+1}$ after time discretizations. I.e., we only evaluate $\psi$ when $t=t_{j}$ ($j=0,1,\ldots,N_{t}$). Here, $\theta^{A}\in\mathbb{R}^{(N_{t}+1)\times d\times d}_{\text{sym}}$ means that each slice of matrix $\theta^{A}_{j}\in\mathbb{R}^{d\times d}$ is symmetric. This parametrization is different from equation 8, but the numerical implementation for the ODE discretization equation 9-equation 13 and loss cost equation 31 remain unchanged, with the vector field defined by
@@
In the LQ example with Gaussian distribution, the solution is indeed of this form; see Bensoussan et al. (2013) Chapter 6 for details.
With such parametrizations, we can compute the residual of the HJB equation through equation 30 as a regularizer. Here, the term $\partial_{t}\psi(t,z_{t})$ in equation 30 could be approximated using the finite difference scheme. In this work, we use the central difference scheme
@@
With this discretization, the numerical residual of the HJB equation at $t_{j}$ (cf. equation 30) is given by
$\mathcal{L}_{\text{HJB}t_{j}}$ $=\dfrac{1}{N_{z}}\sum_{n=1}^{N_{z}}\left[\dfrac{1}{2\Delta t}\left(\psi(t_{j+1},z^{(n)}_{t_{j}};\theta)-\psi(t_{j-1},z^{(n)}_{t_{j}};\theta)\right)+\frac{1}{2}\left|\nabla_{z}\psi(t_{j},z^{(n)}_{t_{j}};\theta)\right|^{2}\right.$ (86)
$\hskip 57.81621pt\left.+\gamma^{2}\left(\operatorname{Tr}\left(H^{(n)}_{t_{j}}\right)+\frac{1}{2}\left|s^{(n)}_{t_{j}}\right|^{2}\right)\right]\,,$
@@
Finally, we can minimize $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cost}}+\lambda\mathcal{L}_{\text{HJB}}$ w.r.t. $\theta$ to solve the MFC problem, where $\lambda>0$ is a weight parameter. We summarize the regularized method in Appendix B.2 Algorithm 2
Regularized algorithm for flow matching of overdamped Langevin dynamics This regularized algorithm could be extended to a more general setting, such as the flow matching problem for an OU process. We recall that the formulation for this problem is presented in Appendix A.3.
@@
Similar to the algorithm presented above, we parametrize $\psi$ through equation 83. Please note that the solution $\psi$ to the MFC problem is given by Corollary 2. So, the vector field is
@@
$\partial_{t}\psi(t,z_{t})+\frac{1}{2}\left|\nabla_{z}\psi(t,z_{t})\right|^{2}+\nabla_{z}\psi(t,z_{t})^{\top}b(t,z_{t})-\gamma\nabla_{z}\cdot b(t,_{t})+\gamma^{2}\left(\operatorname{Tr}(H_{t})+\frac{1}{2}|s_{t}|^{2}\right)=0\,.$ (92)
One can view equation 92 as an analog of equation 30 in the flow matching example.
The time discretization for the HJB equation is the same as equation 85, and the residual of the HJB equation at $t_{j}$ is given by
$\mathcal{L}_{\text{HJB}t_{j}}$ $=\dfrac{1}{N_{z}}\sum_{n=1}^{N_{z}}\left[\dfrac{1}{2\Delta t}\left(\psi(t_{j+1},z^{(n)}_{t_{j}};\theta)-\psi(t_{j-1},z^{(n)}_{t_{j}};\theta)\right)+\frac{1}{2}\left|\nabla_{z}\psi(t_{j},z^{(n)}_{t_{j}};\theta)\right|^{2}\right.$ (93)
$\left.+\nabla_{z}\psi(t_{j},z^{(n)}_{t_{j}};\theta)^{\top}b(t_{j},z^{(n)}_{t_{j}})-\gamma\nabla_{z}\cdot b(t_{j},z^{(n)}_{t_{j}})+\gamma^{2}\left(\operatorname{Tr}\left(H^{(n)}_{t_{j}}\right)+\frac{1}{2}\left|s^{(n)}_{t_{j}}\right|^{2}\right)\right]\,,$
@@
And the total loss is again $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cost}}+\lambda\mathcal{L}_{\text{HJB}}$. We summarize this method in Algorithm 3 Appendix B.2.

### v2

@@ A.1 Proofs for score-based normalizing flow
If we compute the derivative of the MA equation (1) w.r.t. $t$, we obtain
@@
$=\left[\partial_{t}\rho(t,T(t,x))+\nabla_{T}\rho(t,T(t,x))^{\top}f(t,T(t,x))\right.$
$\quad\left.+\rho(t,T(t,x))\nabla_{T}\cdot f(t,T(t,x))\right]\cdot\det\left(\nabla_{x}T(t,x)\right)$
@@
where we have used (45) in the fourth equality. Since $T(t,\cdot)$ is invertible, $\det\left(\nabla_{x}T(t,x)\right)\neq 0$, which implies
@@
$\partial_{t}z_{t}$ $=f(t,z_{t})\,,$
$\partial_{t}l_{t}$ $=-\nabla_{z}\cdot f(t,z_{t})\,,$
$\partial_{t}s_{t}$ $=-\nabla_{z}f(t,z_{t})^{\top}s_{t}-\nabla_{z}(\nabla_{z}\cdot f(t,z_{t}))\,,$
$\partial_{t}H_{t}$ $=-\sum_{i=1}^{d}s_{it}\nabla_{z}^{2}f_{i}(t,z_{t})-\nabla_{z}^{2}(\nabla_{z}\cdot f(t,z_{t}))-H_{t}\nabla_{z}f(t,z_{t})-\nabla_{z}f(t,z_{t})^{\top}H_{t}\,,$
@@
The dynamic for $z_{t}$ (3a) comes from the definition directly:
@@
We next show (4) and (3b).
@@
where we have used the continuity equation (2) in the fourth equality. Consequently,
@@
We next show (3c). We will use the fact frequently that $\nabla_{z}\log\rho(t,z_{t})=\nabla_{z}\rho(t,z_{t})/\rho(t,z_{t})$. We first compute $\dfrac{\mathrm{d}}{\mathrm{d}t}(\nabla_{z}\rho(t,z_{t}))$
$\quad\dfrac{\mathrm{d}}{\mathrm{d}t}\left[\nabla_{z}\rho(t,z_{t})\right]=(\partial_{t}\nabla_{z})\rho(t,z_{t})+\nabla_{z}^{2}\rho(t,z_{t})\,\partial_{t}z_{t}$ (47)
$=\nabla_{z}\partial_{t}\rho(t,z_{t})+\nabla_{z}^{2}\rho(t,z_{t})\,f(t,z_{t})$
@@
where we have used (47) and (4) in the third inequality.
Finally we prove (3d). In order to simplify notation, we will omit $(t,z_{t})$ when there is no confusion and rewrite $\rho(t,z_{t})$ and $f(t,z_{t})$ as $\rho$ and $f$. We also denote $\partial_{i}$ as the partial derivative w.r.t. the $i$-th variable in $z$. We first compute $\dfrac{\mathrm{d}}{\mathrm{d}t}\nabla_{z}^{2}\rho(t,z_{t})$.
@@
$=-\nabla_{z}^{2}[\nabla_{z}\rho^{\top}f+\rho\,\nabla_{z}\cdot f]+\nabla_{z}^{3}\rho\cdot f$
$=-\nabla_{z}^{2}\left[\sum_{i=1}^{d}\partial_{i}\rho\,f_{i}+\rho\,\nabla_{z}\cdot f\right]+\nabla_{z}^{3}\rho\cdot f$
$=-\sum_{i=1}^{d}\left(\nabla_{z}^{2}\partial_{i}\rho\,f_{i}+\partial_{i}\rho\,\nabla_{z}^{2}f_{i}+\nabla_{z}\partial_{i}\rho\,\nabla_{z}f_{i}^{\top}+\nabla_{z}f_{i}\,\nabla_{z}\partial_{i}\rho^{\top}\right)+\nabla_{z}^{3}\rho\cdot f$
$\quad-\left(\nabla_{z}^{2}\rho\,(\nabla_{z}\cdot f)+\rho\,\nabla_{z}^{2}(\nabla_{z}\cdot f)+\nabla_{z}\rho\,\nabla_{z}(\nabla_{z}\cdot f)^{\top}+\nabla_{z}(\nabla_{z}\cdot f)\,\nabla_{z}\rho^{\top}\right)$
@@
where we have used equation (50) in the second equality. Finally, we get
@@
We have used (49) in the second equality and used (51), (50), (4), and (47) in the fourth equality. This finishes the proof. ∎
@@
if equation (53) holds. We first compute
$\quad\dfrac{\mathrm{d}}{\mathrm{d}t}\det(\Sigma(t))=\det(\Sigma(t))\operatorname{Tr}\left(\Sigma(t)^{-1}\,\partial_{t}\Sigma(t)\right)$
$=\det(\Sigma(t))\operatorname{Tr}\left(\Sigma(t)^{-1}\,(A(t)\Sigma(t)+\Sigma(t)A(t)^{\top})\right)$
$=\det(\Sigma(t))\operatorname{Tr}\left(\Sigma(t)^{-1}A(t)\Sigma(t)+A(t)^{\top}\right)=2\det(\Sigma(t))\operatorname{Tr}(A(t))\,,$
@@
$\dfrac{\mathrm{d}}{\mathrm{d}t}\left(\det(\Sigma(t))^{-\frac{1}{2}}\right)$ $=-\frac{1}{2}\det(\Sigma(t))^{-\frac{3}{2}}\,\dfrac{\mathrm{d}}{\mathrm{d}t}\det(\Sigma(t))$ (55)
$=-\det(\Sigma(t))^{-\frac{1}{2}}\,\operatorname{Tr}(A(t))\,.$
@@
Combining (55) and (57), we obtain
@@ A.2 Details for the MFC problems
$\inf_{v}\int_{0}^{t_{\text{end}}}$ $\int_{\mathbb{R}^{d}}\left[L(t,x,v(t,x))+F(t,x,\rho(t,x))\right]\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t$ (60)
@@
Let $L$ be strongly convex in $v$, then solution to the MFC problem (23) is as follows. Consider a function $\psi\colon[0,t_{\textrm{end}}]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that
@@
$\left.\quad+\phi(t,x)\left(\partial_{t}\rho(t,x)+\nabla_{x}\cdot(f(t,x)\rho(t,x))\right)\right]\,\mathrm{d}x\,\mathrm{d}t$ (70)
$\quad+\int_{\mathbb{R}^{d}}G(x,\rho(t_{\text{end}},x))\rho(t_{\text{end}},x)\,\mathrm{d}x$ (71)
@@
Taking the variation of (73) w.r.t. $\rho(t_{\text{end}},\cdot)$, we get $\phi(t_{\text{end}},x)=-\widetilde{G}(x,\rho(t_{\text{end}},x))$. Taking the variation w.r.t. $f$, we obtain
@@
due to (64). Next, we denote $\mathrm{D}_{p}H(t,x,\nabla_{x}\phi(t,x))$ by $\mathrm{D}_{p}H$ for short. We then observe that
@@
Taking the variation of (73) w.r.t. $\rho(\cdot,\cdot)$, we get
@@
$=-\gamma\Delta_{x}\phi(t,x)+L(t,x,\mathrm{D}_{p}H)+\widetilde{F}(t,x,\rho(t,x))$ (80)
$\quad-\partial_{t}\phi(t,x)-\nabla_{x}\phi(t,x)^{\top}\mathrm{D}_{p}H\,,$ (81)
where we have used $\mathrm{D}_{v}L=\nabla_{x}\phi(t,x)$ in the last inequality. Equation (81) is the classical HJB equation corresponding to the MFC problem, which coincide with equation (4.6) in [13] after a sign flip, where our $\mathrm{D}_{p}H$ is their $\hat{v}$. Please note that, throughout our derivation, $\rho(t,x)$ is not a general density, but the density under the optimal control.
@@
where the last equality is due to (75). Therefore, the HJB equation (81) is transformed into
@@
Here, $\mathrm{D}_{p}H$ is short for $\mathrm{D}_{p}H(t,x,\nabla_{x}\psi(t,x)+\gamma\nabla_{x}\log\rho(t,x))$. We plug in the shift (82) in the first equality; we plug in (83) in the second equality. Therefore, we have recovered Proposition 3. ∎
@@
In the LQ problem where $L(t,x,v)=\frac{1}{2}|v|^{2}$ and $F(t,x,\rho)=0$, the solution of the MFC problem (23) is characterized by
@@ A.3 Formulation for flow matching of overdamped Langevin dynamic
For flow matching problem of overdamped Langevin dynamic where $L(t,x,v)=\frac{1}{2}|v-b(t,x)|^{2}$, $F(t,x,\rho)=0$ and $G(x,\rho)=0$, the solution of the MFC problem (23) is characterized by
@@ B.1 HJB regularizer for MFC
In addition to the numerical algorithm in section 4.3, we can add the residual of the HJB equation (27) into the loss function. First, let us consider the HJB regularizer in the LQ Gaussian case, as presented in Corollary 1. Here the Gaussian case means that $\rho_{0}$ is also given as a Gaussian distribution, in which the solution of density function in LQ MFC problem stays in a time dependent Gaussian distribution.
In the LQ scenario, parametrizing the vector field $f$ directly as a neural network complicates the system, as terms like $\partial_{t}\psi$ in (29) are hard to implement. A more feasible approach is to parametrize $\psi$ as a neural network, allowing the vector field $f$ to be computed as $f(t,x)=\nabla_{x}\psi(t,x)$.
This parametrization, however, requires two additional orders of auto-differentiations compared to the original algorithm. First, if $\psi$ is parametrized as a neural network as in (8), then $f(t,x)=\nabla_{x}\psi(t,x)$ must be obtained through auto-differentiation. Second, the term $\Delta_{x}\rho(t,x)$ in (29) requires computing the dynamic for $H_{t}$ through (13), which requires another order of derivative. This term $H_{t}$ is not required in the original Algorithm 1. These two additional differentiations could potentially complicate the optimization landscape.
@@
where $A(t):[0,t_{\text{end}}]\to\mathbb{R}^{d\times d}$, $B(t):[0,t_{\text{end}}]\to\mathbb{R}^{d}$, and $C(t):[0,t_{\text{end}}]\to\mathbb{R}$. These three functions can be further parametrized into trainable variables $\theta=[\theta^{A},\theta^{B},\theta^{C}]$ with $\theta^{A}\in\mathbb{R}^{(N_{t}+1)\times d\times d}_{\text{sym}}$, $\theta^{B}\in\mathbb{R}^{(N_{t}+1)\times d}$, and $\theta^{C}\in\mathbb{R}^{N_{t}+1}$ after time discretizations. I.e., we only evaluate $\psi$ when $t=t_{j}$ ($j=0,1,\ldots,N_{t}$). Here, $\theta^{A}\in\mathbb{R}^{(N_{t}+1)\times d\times d}_{\text{sym}}$ means that each slice of matrix $\theta^{A}_{j}\in\mathbb{R}^{d\times d}$ is symmetric. This parametrization is different from (8), but the numerical implementation for the ODE discretization (9)-(13) and loss cost (31) remain unchanged, with the vector field defined by
@@
In the LQ example with Gaussian distribution, the solution is indeed of this form; see [13] Chapter 6 for details.
With such parametrizations, we can compute the residual of the HJB equation through (30) as a regularizer. Here, the term $\partial_{t}\psi(t,z_{t})$ in (30) could be approximated using the finite difference scheme. In this work, we use the central difference scheme
@@
With this discretization, the numerical residual of the HJB equation at $t_{j}$ (cf. (30)) is given by
$\mathcal{L}_{\text{HJB}t_{j}}$ $=\dfrac{1}{N_{z}}\sum_{n=1}^{N_{z}}\left[\dfrac{1}{2\Delta t}\left(\psi(t_{j+1},z^{(n)}_{t_{j}};\theta)-\psi(t_{j-1},z^{(n)}_{t_{j}};\theta)\right)\right.$ (92)
$\hskip 28.90755pt\left.+\frac{1}{2}\left|\nabla_{z}\psi(t_{j},z^{(n)}_{t_{j}};\theta)\right|^{2}+\gamma^{2}\left(\operatorname{Tr}\left(H^{(n)}_{t_{j}}\right)+\frac{1}{2}\left|s^{(n)}_{t_{j}}\right|^{2}\right)\right]\,,$
@@
Finally, we can minimize $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cost}}+\lambda\mathcal{L}_{\text{HJB}}$ w.r.t. $\theta$ to solve the MFC problem, where $\lambda>0$ is a weight parameter. We summarize the regularized method in B.2 Algorithm 2.
Regularized algorithm for flow matching of overdamped Langevin dynamics. This regularized algorithm could be extended to a more general setting, such as the flow matching problem for an OU process. The formulation for this problem is presented in A.3.
@@
Similar to the algorithm presented above, we parametrize $\psi$ through (89). Please note that the solution $\psi$ to the MFC problem is given by Corollary 2. So, the vector field is
@@
$\partial_{t}\psi(t,z_{t})+\frac{1}{2}\left|\nabla_{z}\psi(t,z_{t})\right|^{2}+\nabla_{z}\psi(t,z_{t})^{\top}b(t,z_{t})$ (98)
$-\gamma\nabla_{z}\cdot b(t,_{t})+\gamma^{2}\left(\operatorname{Tr}(H_{t})+\frac{1}{2}|s_{t}|^{2}\right)=0\,.$
One can view (98) as an analog of (30) in the flow matching example.
The time discretization for the HJB equation is the same as (91), and the residual of the HJB equation at $t_{j}$ is given by
$\mathcal{L}_{\text{HJB}t_{j}}$ $=\dfrac{1}{N_{z}}\sum_{n=1}^{N_{z}}\left[\dfrac{1}{2\Delta t}\left(\psi(t_{j+1},z^{(n)}_{t_{j}};\theta)-\psi(t_{j-1},z^{(n)}_{t_{j}};\theta)\right)\right.$ (99)
$+\frac{1}{2}\left|\nabla_{z}\psi(t_{j},z^{(n)}_{t_{j}};\theta)\right|^{2}+\nabla_{z}\psi(t_{j},z^{(n)}_{t_{j}};\theta)^{\top}b(t_{j},z^{(n)}_{t_{j}})$
$\left.-\gamma\nabla_{z}\cdot b(t_{j},z^{(n)}_{t_{j}})+\gamma^{2}\left(\operatorname{Tr}\left(H^{(n)}_{t_{j}}\right)+\frac{1}{2}\left|s^{(n)}_{t_{j}}\right|^{2}\right)\right]\,,$
@@
And the total loss is again $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cost}}+\lambda\mathcal{L}_{\text{HJB}}$. We summarize this method in Algorithm 3 B.2.